OpenStudy (anonymous):

Suppose each face of a 3cm X 5cm X 4cm block of wood is shaven down by x cm. If the shaved block has half the original volume, find the value of x to the nearest hundredth.

OpenStudy (paxpolaris):

what's the original volume?

OpenStudy (anonymous):

well (3)(4)(5)=60 therefore 60

OpenStudy (paxpolaris):

well then if you subtract x from each side you get: 3-x, 4-x, 5-x

OpenStudy (anonymous):

would i multiply them together and set it up to 30?

OpenStudy (paxpolaris):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

@PaxPolaris can I do something like: the shaved one has volume 30 so, the taken out part has volume is 30 also, so that x*x*x =30 and get x by that way?

OpenStudy (anonymous):

That's what I was thinking but I wasn't sure

OpenStudy (anonymous):

Also once I set the equation up, would I distribute and solve or is there a more efficient way?

OpenStudy (paxpolaris):

@OOOPS no x^3 =30 won't work

OpenStudy (anonymous):

can i graph this equation on my calculator and find the maximum of the x value @PaxPolaris

OpenStudy (anonymous):

oh yea, I see my mistake. I am sorry. You are right. @PaxPolaris

OpenStudy (paxpolaris):

yes, i think you have to graph to get the answer find the point of intersection of \( y= (3-x)(4-x)(5-x)\) and \(y=30\) or see where \(y= (3-x)(4-x)(5-x)-30\) meets the x-axis

OpenStudy (anonymous):

I got x=0.78

OpenStudy (paxpolaris):

yep . :)

OpenStudy (anonymous):

Thank you so much!